OpenStudy (anonymous):

A 950 gallon swimming pool is being drained. After seven hours, 500 gallons have been drained. If the rate of drainage continues, how long will it take to drain the pool?

OpenStudy (anonymous):

500/7

OpenStudy (anonymous):

71.42857143

OpenStudy (anonymous):

^gallons per hour

OpenStudy (anonymous):

950/ 71.42857143

hero (hero):

x = drainage in hours y = gallons remaining (x,y) = relationship between drainage and gallons You are given (0, 950) and (7, 500) Find the slope using \(m = \dfrac{y_2 - y_1}{x_2 - x_1}\) Then use \(y = mx + b\) to write the equation of the line.

hero (hero):

Once you find the equation of the line, let y = 0, then solve for x.